\begin{thebibliography}{1}

\bibitem{Yao:1993:QCircuits}
Andrew Chi-Chih Yao.
\newblock Quantum circuit complexity.
\newblock In {\em Proceedings of Thirty-fourth IEEE Symposium on Foundations of
  Computer Science}, pages 352--361, Palo Alto, California, USA, 1993. IEEE
  Computer Society.

\bibitem{Jain:1992:verification}
Jawahar Jain, Jacob~A. Abraham, James Bitner, and Donald~S. Fussell.
\newblock Probabilistic verification of boolean functions.
\newblock {\em Formal Methods in System Design}, 1:61--115, 1992.

\bibitem{Agrawal:1998:characteristic-polynomials}
Vishwani Agrawal, David Lee, and Henryk Wozniakowski.
\newblock Numerical computation of characteristic polynomials of boolean
  functions and its applications.
\newblock {\em Numerical Algorithms}, 17:261--278, 1998.

\bibitem{Buhrman:2001:Fingerprinting}
Harry Buhrman, Richard Cleve, John Watrous, and Ronald de~Wolf.
\newblock Quantum fingerprinting.
\newblock {\em Phys. Rev. Lett.}, 87(16):167902, Sep 2001.

\bibitem{Wolf:2001:PhD}
Ronald de~Wolf.
\newblock {\em Quantum Computing and Communication Complexity}.
\newblock PhD thesis, University of Amsterdam, 2001.

\bibitem{ablayev-vasiliev:2009:EPTCS}
Farid Ablayev and Alexander Vasiliev.
\newblock Algorithms for quantum branching programs based on fingerprinting.
\newblock {\em Electronic Proceedings in Theoretical Computer Science},
  9:1--11, 2009.

\bibitem{Ablayev-Vasiliev:2013:Quantum-Hashing}
Farid Ablayev and Alexander Vasiliev.
\newblock Quantum hashing.
\newblock Technical Report arXiv:1310.4922 [quant-ph], Cornell University
  Library, Oct 2013.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsfonts}
\usepackage[margin=2cm,nohead]{geometry}

\usepackage[normalem]{ulem}

\newcommand{\comment}[1]{}

\newcommand{\braket}[2]{{\langle {#1}\!\mid\!{#2} \rangle}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{definition}{Definition}[section]
\newtheorem{lemma}{Lemma}[section]
\newtheorem{corollary}{Corollary}[section]

\newtheorem{remark}{Remark}[section]

\newcommand{\mat}{\left( \!\! \begin{array}{cc}}
\newcommand{\rix}{\end{array} \!\! \right)}

\newcommand{\Endproof}{\hfill$\Box$\\}

\usepackage{color}
\def\bR{\begin{color}{red}}
\def\bB{\begin{color}{blue}}
\def\bM{\begin{color}{magenta}}
\def\bC{\begin{color}{cyan}}
\def\bW{\begin{color}{white}}
\def\bBl{\begin{color}{black}}
\def\bG{\begin{color}{green}}
\def\bY{\begin{color}{yellow}}
\def\ec{\end{color}\ }

\usepackage{xy}
\xyoption{matrix}
\xyoption{frame}
\xyoption{arrow}
\xyoption{arc}

\usepackage{ifpdf}
\ifpdf
\else
\PackageWarningNoLine{Qcircuit}{Qcircuit is loading in Postscript mode.  The Xy-pic options ps and dvips will be loaded.  If you wish to use other Postscript drivers for Xy-pic, you must modify the code in Qcircuit.tex}
\xyoption{ps}
\xyoption{dvips}
\fi

\entrymodifiers={!C\entrybox}

\newcommand{\ket}[1]{{\left\vert{#1}\right\rangle}}

\begin{document}

\title{Quantum Communications Based on Quantum Hashing}

\author{Alexander Vasiliev\thanks{Kazan Federal University}}

\date{}


\maketitle

\begin{abstract}

Based on the polynomial presentation of Boolean functions and
quantum hashing technique we present a method for computing Boolean
functions in the quantum one-way communication model. Some of the
results are also true in a more restricted Simultaneous Message
Passing model.
\end{abstract}



\section{Introduction}

While a large-scale fully functional quantum computer remain a
theoretical model, quantum communications are extensively
implemented and may soon enter our everyday life. That is why the
study of different quantum communication models may add value to
this technology. However, in absence of long-term quantum memory and
quite small coherence time of quantum states we should consider
restricted versions of quantum communication models in the first
place. In particular, such models those considered here: the one-way
quantum communication model and the simultaneous message passing
model \cite{Yao:1993:QCircuits} with no shared resources.

From the complexity theoretic viewpoint such a strong restrictions
on a computational model give way to a variety of techniques for
proving lower bounds on the complexity in this model. Sometimes it
even allows to prove the asymptotic optimality of the protocols.

Our approach relies on the polynomial presentation of Boolean
functions which has proven its usefulness in a number of papers
\cite{Jain:1992:verification},\cite{Agrawal:1998:characteristic-polynomials},
\cite{Buhrman:2001:Fingerprinting},\cite{Wolf:2001:PhD}. However,
here we use a slightly different type of polynomial presentation
proposed in \cite{ablayev-vasiliev:2009:EPTCS}.

Another component of our approach is quantum hashing
\cite{Ablayev-Vasiliev:2013:Quantum-Hashing}, which transforms a
classical input into quantum superposition. Here, hashing is used to
reduce the amount of data transferred between communicating parties,
just like it has been done by means of quantum fingerprinting in
\cite{Buhrman:2001:Fingerprinting}. Actually, quantum fingerprinting
is also quantum hashing in terms of
\cite{Ablayev-Vasiliev:2013:Quantum-Hashing}.

Finally, the main construction proves the existence of an effective
quantum communication protocols for the class of functions with
specific polynomial presentations. Several known Boolean functions
from this class are exposed.

\section{Preliminaries}

At the core of our approach lies the polynomial presentation of
Boolean functions proposed in \cite{ablayev-vasiliev:2009:EPTCS}. We
recall some of the definitions here.

\paragraph{Polynomial presentation of
Boolean functions.}
\begin{definition}\label{polynomial-definition}
We call a polynomial $g(x_1,\dots,x_n)$ over the ring ${\mathbb
Z}_m$ a characteristic polynomial of a Boolean function
$f(x_1,\ldots,x_n)$ and denote it $g_f$
%
when for all $\sigma\in\{0,1\}^n$ 
$g_f(\sigma)=0$ iff $f(\sigma)=1$.
\end{definition}

It was also shown that such a polynomial always exists (but is not
unique).

\begin{lemma}\label{Existence-Of-A-Characteristic-Polynomial}
For any Boolean function $f$ of $n$ variables there exists a
characteristic polynomial $g_f$ over ${\mathbb Z}_{2^n}$.
\end{lemma}

We have used this presentation to test a single property of the
input encoded by a characteristic polynomial. Using the same ideas
we can test the conjunction of several conditions encoded by a group
of characteristic polynomials which we call a \emph{characteristic}
of a function.

\begin{definition}
We call a set $\chi_f$ of polynomials over $\mathbb{Z}_m$ a
\emph{characteristic} of a Boolean function $f$ if for all
polynomials $g\in\chi_f$ and any $\sigma\in\{0,1\}^n$ it holds that
$g(\sigma)=0$ iff $\sigma\in f^{-1}(1)$.
\end{definition}

We say that a characteristic is \emph{linear} if all of its
polynomials are linear. In \cite{ablayev-vasiliev:2009:EPTCS} we
have shown that Boolean functions with linear characteristics of
logarithmic size can be efficiently computed in the quantum OBDD
model.

\paragraph{Quantum Hashing}

We recall a definition of quantum hashing function from
\cite{Ablayev-Vasiliev:2013:Quantum-Hashing}.

Let $N=2^n$. Let $K=\{k_i: k_i\in\{0,\ldots,N-1\}\}$ and $d=|K|$. We define a classical-quantum function
\[ h_K : \{0,1\}^n \to ({\mathbb C}^2)^{\otimes (d+1) } \]
as follows. For a message $M\in\{0,1\}^n$ we let
%
\[
\ket{h_K(M)} =
\frac{1}{\sqrt{d}}\sum\limits_{i=1}^d\ket{i}
\left(\cos\frac{2\pi k_i
M}{N}\ket{0}+\sin\frac{2\pi k_i M}{N}\ket{1}\right).
\]
The set $K$ is used to bound the probability of collisions of was
proven to exist for any $\delta\in(0,1)$ and $N$ with $|K|=O((\log
N)/\delta)$ in \cite{Ablayev-Vasiliev:2013:Quantum-Hashing}.

\section{Quantum Communication Protocols Based on Quantum Hashing}

The quantum hashing defined above can be used for constructing effective protocols
in the quantum communication model defined by Yao in \cite{Yao:1993:QCircuits}.

Here we consider a one-sided restriction of this model, where Alice makes her computations,
sends some information to Bob, who computes his part of the protocol and outputs the result.
The complexity of such a protocol is the number of qubits sent to Bob.

Let $f(x_1,\ldots,x_{n_1}, y_1,\ldots, y_{n_2})$ be a Boolean
function of $n=n_1+n_2$ variables, i.e.
$$f:\{0,1\}^{n_1} \times \{0,1\}^{n_2}  \to \{ 0,1\}$$.
Alice gets the sequence of values
$\sigma=\sigma_1\ldots\sigma_{n_1}$ of the first $n_1$ variables, and Bob
gets $\gamma=\gamma_1\ldots\gamma_{n_2}$  -- the values of the last
$n_2$ variables.

To compute $f$ we exploit its polynomial presentation, proposed in
\cite{ablayev-vasiliev:2009:EPTCS}. Namely, we call a polynomial
$g_f(x_1,\dots,x_n)$ over the ring ${\mathbb Z}_N$ a characteristic polynomial
of a Boolean function $f(x_1,\ldots,x_n)$ if for all $\sigma\in\{0,1\}^n$
$g_f(\sigma)=0$ iff $f(\sigma)=1$. We have shown that such a polynomial over
${\mathbb Z}_{2^n}$ exists for any Boolean function of $n$ variables, but is
not unique.

In the communication scenario the input is split between parties, and a polynomial for $f$ should also be decomposed.
For the quantum hashing technique proposed we decompose this polynomial into the sum of two polynomials, one for each of the
communicating parties.


\begin{theorem}\label{quantum-communication-computation}
Let $f(x_1,\ldots,x_{n_1}, y_1,\ldots, y_{n_2})$ be a Boolean function of $n=n_1+n_2$ variables.
Let $g$ be a characteristic polynomial for $f$ over a ring
${\mathbb Z}_N$. If  $g$ can be decomposed into
$$g(x_1,\ldots,x_{n_1}, y_1,\ldots, y_{n_2})=g_1(x_1,\ldots,x_{n_1})+g_2(y_1,\ldots, y_{n_2}),$$
then for arbitrary $\delta >0$ $f$ can be computed by a one-way quantum communication protocol
with $O(\log\log{N}+\log(1/\delta))$ qubits of communication.
\end{theorem}
{\em Proof.} For the proof we describe the following quantum one-way communication protocol.

The communicating parties given an input $(\sigma, \gamma)$ want to know whether $f(\sigma, \gamma)=1$ or
not. This is the same as asking whether $g(\sigma, \gamma)=0$, or, equivalently, whether
$g_1(x_1,\ldots,x_{n_1})=-g_2(y_1,\ldots, y_{n_2})$. And this equality is exactly what the
protocol would check using quantum hashing technique, i.e. it will compare quantum hashes of those
values.

More formally, the following describes a one-way protocol of computing $f$
in the quantum communication setting using $\delta$-collision resistant quantum
hashing for some $\delta\in(0,1)$.

1. Alice, depending on her input $\sigma=\sigma_1\ldots \sigma_{n_1}$, creates
a quantum hash for the value $g_1(\sigma)$
\[
\ket{h_{g_1(\sigma)}} =
\frac{1}{\sqrt{d}}\sum\limits_{j=1}^d\ket{j}
\left(\cos\frac{2\pi k_j
g_1(\sigma)}{N}\ket{0}+\sin\frac{2\pi k_j g_1(\sigma)}{N}\ket{1}\right)
\]
and sends it to Bob.

2. Given $\ket{h_{g_1(\sigma)}}$ and his input $\gamma=\gamma_1\ldots\gamma_{n_2}$ Bob
creates a quantum hash for the value $-g_2(\gamma)$
\[\begin{array}{rcl}
\ket{h_{-g_2(\gamma)}} & = &
\frac{1}{\sqrt{d}}\sum\limits_{i=1}^d\ket{i}\left(\cos\frac{2\pi k_i
(-g_2(\gamma))}{N}\ket{0}+\sin\frac{2\pi k_i (-g_2(\gamma))}{N}\ket{1}\right)
\end{array}\]

3. Bob compares $\ket{h_{g_1(\sigma)}}$ and $\ket{h_{-g_2(\gamma)}}$ using the
SWAP-test. So, Bob obtains the result $\ket{h_{u}}=\ket{h_{v}}$, if the measurement of the first
qubit gives $\ket{0}$, which happens with probability
$\frac{1}{2}\left(1+|\braket{h_{u}}{h_{v}}|^2\right)$.

4. Bob outputs the result of computations. He says $f(\sigma,\gamma) = 1$ if $\ket{h_{g_1(\sigma)}}$ = $\ket{h_{-g_2(\gamma)}}$
and $f(\sigma,\gamma) = 0$ otherwise.

If the value of $f(\sigma,\gamma)$ was 1, then Bob outputs 1 with
certainty. If $f(\sigma,\gamma)$ was 0, then by $\delta$-resistance
property
$\left|\braket{h_{g_1(\sigma)}}{h_{-g_2(\gamma)}}\right|<\delta$,
and the probability of erroneously outputting 1 is bounded by $1/2 +
\delta^2/2$.

The communication complexity in this case is bounded by the size of the quantum
hash passed from Alice to Bob, which is $\log{d}+1$ $=$
$O\left(\log\log{N}+\log(1/\delta)\right)$ qubits. \Endproof

We now recall our assumption that characteristic polynomial can be decomposed into the
sum of polynomials over independent sets of variables. The simplest case of such polynomials
are linear polynomials and we have exposed in \cite{ablayev-vasiliev:2009:EPTCS} several examples
of natural Boolean functions that have linear characteristic polynomials. Among them there is
an \emph{Equality test}, which is a frequently considered in the study of communication complexity.
The corresponding Boolean function has the following linear characteristic polynomial over ${\mathbb Z}_{2^n}$
$$g_{EQ}(x_1,\ldots,x_{n}, y_1,\ldots, y_{n})=\sum\limits_{i=1}^nx_i2^{i-1} - \sum\limits_{i=1}^ny_i2^{i-1},$$
and thus can be computed by the $O(\log{n})$-qubit quantum
communication protocol.

Here are some more Boolean functions with linear characteristic
polynomials, which are thus effectively computable in the SMP model.

\paragraph{$MOD_m$} The function $MOD_m$ tests whether the number of
$1$'s in the input is $0$ modulo $m$. The linear polynomial over
$\mathbb{Z}_m$ for this function is
\[\sum\limits_{i=1}^nx_i.\]

\paragraph{$MOD'_m$} This function is the same as $MOD_m$, but the input
is treated as binary number. Thus, the linear polynomial is
\[\sum\limits_{i=1}^nx_i2^{i-1}.\]


\paragraph{$Palindrome_n(x_1,\ldots,x_n)$} This function tests the symmetry of the
input, i.e. whether $x_1 x_2\ldots x_{\lfloor n/2\rfloor}$ = $x_n
x_{n-1}\ldots x_{\lceil n/2\rceil+1}$ or not. The polynomial over
$\mathbb{Z}_{2^{\lfloor n/2\rfloor}}$ is
\[\sum\limits_{i=1}^{\lfloor n/2\rfloor}x_i2^{i-1}- \sum\limits_{i=\lceil n/2\rceil}^nx_i2^{n-i}.\]

\paragraph{$PERM_n$} The \emph{Permutation Matrix} test function ($PERM_n$) is defined on
$n^2$ variables $x_{i j}$ ($1\leq i,j\leq n$). It tests whether the
input matrix contains exactly one 1 in each row and each column.
Here is a polynomial over $\mathbb{Z}_{(n+1)^{2n}}$
\[\sum\limits_{i=1}^n\sum\limits_{j=1}^nx_{ij}\left((n+1)^{i-1}+(n+1)^{n+j-1}\right)
            - \sum\limits_{i=1}^{2n}(n+1)^{i-1}.\]

\begin{remark}
By inspecting the proposed communication protocol one can note,
that it is still valid in a more restrictive setting of \emph{simultaneous
message passing model}, and the Theorem \ref{quantum-communication-computation} can be
restated and proved for this model as well.
\end{remark}





Now, if for some Boolean function $f$ there is no characteristic polynomial, that can be
decomposed as shown earlier, we use the following decomposition
$$g(x_1,\ldots,x_{n_1}, y_1,\ldots, y_{n_2})=g_1(x_1,\ldots,x_{n_1})+g_2(x_{i_1},\ldots, x_{i_k},y_1,\ldots, y_{n_2}).$$
Such a decomposition always exists, since we can $k=n_1$, $g_1\equiv0$ and
$g_2\equiv g$.

Then the following result holds, which generalizes Theorem \ref{quantum-communication-computation}.

\begin{theorem}\label{general-quantum-communication-computation}
For arbitrary $\delta >0$ $f$ can be computed by a one-way quantum
communication protocol with $O(k+\log\log{N}+\log(1/\delta))$ qubits
of communication.
\end{theorem}

{\em Proof.}
The protocol is almost the same as the one from from Theorem \ref{quantum-communication-computation},
but Alice sends a hash plus $k$ qubits containing the values of
$x_{i_1},\ldots, x_{i_k}$, and Bob use them to construct his own hash.
The protocol now requires
$O\left(k+\log\log{N}+\log(1/\delta)\right)$ qubits of communication.
\Endproof

\begin{corollary}
If $N=2^{n^{O(1)}}$ (which is the most usual case) and $k=O(\log{n})$ the
described protocol would require $O(\log{n})$
qubits of communication, which is
exponentially better than just sending all of the input from Alice to Bob.
\end{corollary}

\begin{remark}
However, for arbitrary Boolean function the bound is still no better than trivial $O(n)$,
since in general $k=O(n)$.
\end{remark}

\section{General Approach}



In a more general approach we consider a characteristic $\chi_f^m$
for some Boolean function $f$ on $n_1+n_2$ variables.

0. We start with fixing some $\delta\in(0,1)$ and two sets
$G=\{g_1,\ldots, g_l\}$ and $R=\{r_1,\ldots, r_l\}$ of polynomials
over the ring ${\mathbb Z}_{m}$, such that the set
$\chi_f^m=\{g_1+r_1,\ldots, g_l+r_l\}$ is a characteristic of $f$
over ${\mathbb Z}_{m}$. Here we assume that polynomials from $G$
depend only on $x_1,\ldots,x_{n_1}$, and those from $R$ -- depend
not only on $y_1,\ldots, y_{n_2}$, but also on $x_{i_1},\ldots,
x_{i_k}$.

1. The protocol starts when $A$ receives an input
$\sigma=\sigma_1\ldots \sigma_{n_1}$, computes $l$ quantum hashes
for the values $g_1(\sigma)$, \ldots, $g_l(\sigma)$, and sends them
to $B$ along with the values $x_{i_1},\ldots, x_{i_k}$.

2. $B$ receives his part of the input
$\gamma=\gamma_1\ldots\gamma_{n_2}$, $l$ quantum hashes and values
$x_{i_1},\ldots, x_{i_k}$. Then he computes $l$ hashes for
$-r_1(\gamma)$, \ldots, $-r_l(\gamma)$ and performs pairwise
comparison using SWAP-test.

3. $B$ outputs $1$ iff all the pairs have passed the test. When
$f(\sigma,\gamma) = 1$, this protocol would always lead to correct
results. But if $f(\sigma,\gamma) = 0$, then for at least one
$i\in\{1,\ldots,l\}$ $g_i(\sigma)\neq-r_i(\gamma)$, then by
$\delta$-resistance property the probability of erroneously
outputting 1 is bounded $\delta/2 + 1/2$.

Thus, the complexity of communication protocols based on quantum
hashing and general characteristic polynomial presentation of
Boolean functions is $O\left(k+|\chi_f^m|\cdot\log\log{m}\right)$.
Whenever $k=O(\log{n})$, $|\chi_f|=O(1)$, and $m=2^{n^{O(1)}}$, the
complexity of such protocol would be $O(\log{n})$
An example of such function is a Boolean version of Hidden Subgroup
Problem, considered in \cite{ablayev-vasiliev:2009:EPTCS}, which has
a characteristic over ${\mathbb Z}_{2^n}$, consisting of two
polynomials.

\section*{Acknowledgements.}

Research was supported by the Russian Fund for Basic Research (under the grants
11-07-00465, 12-01-31216).

\bibliography{references}

\end{document}